OpenStudy (jessiegonzales):

MEDAL!!! Find the surface area of the sphere with the given dimension. Leave your answer in terms of pi. 18. radius of 70 m A.4,900pi m^2 B.19,600pi m^2 C.9,800pi m^2 D.14,700pi m^2

OpenStudy (anonymous):

4*pi*70*70 :)

OpenStudy (jessiegonzales):

wait what? i just multiply the four numbers

OpenStudy (anonymous):

yes :)

OpenStudy (jessiegonzales):

4*3.14=12.56 12.56*70=879.2 879.2*70=61,544

OpenStudy (anonymous):

your choice is in terms of pi, so dont multiply pi

OpenStudy (fibonaccichick666):

do you know the formula for surface area of a sphere?

OpenStudy (jessiegonzales):

19,600pi m^2 thanks

OpenStudy (anonymous):

:)